# Mind the Style Gap: Meta-Evaluation of Style and Attribute Transfer Metrics

### v3

@@ Abstract
Large language models (LLMs) make it easy to rewrite a text in any style – e.g. to make it more polite, persuasive, or more positive – but evaluation thereof is not straightforward. A challenge lies in measuring content preservation: that content not attributable to style change is retained. This paper presents a large meta-evaluation of metrics for evaluating style and attribute transfer, focusing on content preservation. We find that meta-evaluation studies on existing datasets lead to misleading conclusions about the suitability of metrics for content preservation. Widely used metrics show a high correlation with human judgments despite being deemed unsuitable for the task – because they do not abstract from style changes when evaluating content preservation. We show that the overly high correlations with human judgment stem from the nature of the test data. To address this issue, we introduce a new, challenging test set specifically designed for evaluating content preservation metrics for style transfer. Using this dataset, we demonstrate that suitable metrics for content preservation for style transfer indeed are style-aware. To support efficient evaluation, we propose a new style-aware method that utilises small language models, obtaining a higher alignment with human judgements than prompting a model of a similar size as an autorater.
@@ 1 Introduction
Figure 1: Sample from our new test set with human annotations. S: style strength, C: content preservation. Metrics for content preservation: Bold indicates the highest rated output, and the checkmark is correct.
@@
Our aggregated results (Figure 2) demonstrate that similarity-based metrics for content preservation show low to negative correlations with human judgment on our test set (last column), but exhibit overly high correlations to human judgment on existing data sets (first two columns). Style-aware approaches show positive correlations to human judgments on all data types (all columns).
Figure 2: Average Spearman correlation between metrics and human judgments for content preservation, grouped by whether test data originates from system-generated output (total of 5 sources), human-written (total of 3 sources) or our more challenging constructed test set. Which metric is deemed best changes depending on the test data, and only our test set correctly finds similarity-based metrics to be unsuitable.
@@ 2 Background and Related Work
Evaluation. Traditionally, style transfer systems have been evaluated in three dimensions: 1) style strength (did the style successfully shift?); 2) content preservation (is the content otherwise the same?); 3) fluency (is the rewrite fluent and grammatically correct?) Jin et al. (2022), (fluency is not covered in this work). Some work on style transfer systems includes small-scale human evaluations of the systems, but papers heavily rely on automatic metrics for evaluation Ostheimer et al. (2023):
Style. The de facto evaluation approach, also in recent work, uses a classifier, e.g. RoBERTa, trained on a specific style Lai et al. (2024); Hallinan et al. (2023); Han et al. (2024); Mukherjee et al. (2024); Liu et al. (2024); Luo et al. (2023); Zeng et al. (2024). However, this requires training data for the specific style, limiting its scope and scale, and often resorts to a binary classifier, not capturing that a style can shift to differing degrees Briakou et al. (2021a). In this paper, we focus on evaluating arbitrary styles without trained classifiers – and output scores as opposed to binary assessments of style change success.
@@
A proposed solution to this criticism is to make metrics for content preservation style-aware. Mir et al. (2019) proposes to use a lexicon of style words to mask the sentence in a sentiment task. In a follow-up study, Yu et al. (2021) extract a style lexicon automatically. More recent work prompts an LLM for evaluation (LLM as an autorater) Zeng et al. (2024); Lai et al. (2023); Ostheimer et al. (2024). For other natural language generation tasks, LLM-as-a-judge approaches are deemed better than similarity-based metrics when tasks are semantically nuanced Li et al. (2024).
Still, semantic similarity metrics often show reasonable correlation with human judgment on content preservation in style transfer. Prior meta-evaluation efforts (Lai et al., 2023; Ostheimer et al., 2024) show that using LLMs for evaluating content preservation is only on par rather than superior to using semantic similarity-based metrics.
@@ 3 Methodology: Test Set
Human annotators rate style shift on a 1–5 scale (5 = very good). Both outputs are seen as successfully shifting to a more polite tone, but Output 2 scores poorly on content preservation. Yet, similarity-based metrics wrongly rate Output 2 as better at preserving content than Output 1, while style-aware metrics correctly assign it a lower score. (Metrics are formally introduced in Section 5.)
@@
We aim to construct data where all output sentences succeed on style transfer, but for each source sentence, only one output preserves the content well, creating variation in content preservation quality. Our constructed test data covers six style/attribute tasks and consists of 500 samples, 100 of which are manually created and the remaining synthetically generated for scale:
Manually generated. We construct 100 transfer samples from a source sentence by manually creating two output sentences as described. We create: Task 1) on sentiment (positive/negative), using headlines from Wiki News222(Wikinews.org) as source sentences with minor modifications. Task 2) on detoxifying, using toxic sentences from Logacheva et al. (2022b). We supply each sample with a gold-truth reference sentence to enable a small evaluation of reference-based methods.
Synthetically generated. We use GPT-4o-mini333openai.com to create style transfer pairs in a multi-step process:
@@
We construct the synthetic test set with four different style transfer tasks: i) a headline to be more catchy; ii) an impolite sentence from an email to be polite; iii) a persuasive request to be more persuasive; and iv) a sentence with informal language with grammatical mistakes and internet slang to be formal. Prompt details are included in App. B.
@@ 3.1 Human Annotations
We multi-annotate our samples and obtain a good level of agreement: Three workers annotate each sample in batches, grouped by style task, with a total of five different workers recruited via a crowdsourcing platform. Annotators rate, on a 5-point Likert scale, how well the style or attribute change is achieved, and how well the meaning/content unrelated to the style/attribute change is preserved, following previous work Mir et al. (2019); Ziegenbein et al. (2024). We use the scale
@@
We achieve a high level of human agreement on content preservation with a Krippendorff’s Alpha Krippendorff (2011) of $0.768$, and a good level of success in style transfer, with an average score on the 5-point Likert scale of $4.27$. Detailed results per task in Table 1, where we report the percentage of samples that at least obtain an average rating of 3 (fair). Most sentences successfully transfer style. Details on annotation guideline, setup and payment are provided in App. C.
@@
catchy sy 100 0.806 90
polite sy 100 0.645 100
persuasive sy 100 0.80 88
formal sy 100 0.817 99
Table 1: Our test set. Constructed ma: manually, sy: synthetic. C: content preservation, S: style strength
@@ 4 Methodology: LogProb
We propose a novel approach for evaluating style transfer with the following properties: 1) no training data required on style; 2) no need for gold truth (/reference) sentences; 3) improved feasibility in the form of model size (e.g. 1B,3B) with improved performance compared to our baseline of prompting an LLM-as-Autorater. For content preservation, we additionally aim for: 4) accounting for changes in target style – ensuring theoretical suitability as discussed in Section 2. For style strength, beyond properties 1,2 and 3, we aim for: 5) scoring on a range rather than a binary score – motivated by the fact that styles can shift to various degrees as discussed in Section 2 and recommended by Briakou et al. (2021a).
@@
and $t^{s}$,$t^{pa}$,$t^{r}$ being instructions on rewrite with respect to style, paraphrase and repeat.
@@
$S=\frac{1}{n}\sum_{i=1}^{n}\left(p_{i}^{t^{s}}-\max\left(p_{i}^{t^{t}a},p_{i}^{t^{r}}\right)\right)$ (2)
We deploy three backbone model sizes—1B,3B and 8B parameters—using Llama 3 Instruct models Dubey et al. (2024)444 meta-llama/Llama-3.2-3B-Instruct,meta-llama/Llama-3.2-1B-Instruct and meta-llama/Llama-3.1-8B-Instruct. We deploy different instructions as part of the context as per Table 2. Further details in App. D.
@@ 5 Benchmarking
    Semantic similarity: token-based BERTScore Zhang et al. (2019); Cosine similarity between sentence embeddings Feng et al. (2022); learned BERT-based BLEURT Sellam et al. (2020). COMET Li et al. (2024), which uses both source and references to evaluate.
@@ 6 Results
Meta-evaluation is highly impacted by the type of test samples. Overall metric performance (correlation with human judgements) varies greatly depending on whether the human-rated outputs are system-generated, human-written, or from our newly constructed test set (Fig. 2). On system-generated data (Table 3): Many metrics show relatively high correlations across tasks. For all similarity-based metrics, the versions using source sentences [S-] result in higher correlations to human ratings than the versions using references [R-]. The best metrics overall are S-BertScore and S-BLEURT (avg+avg rank). On gold reference data (Table 3): Here, LLM-as-Autorater 70B performs best (avg+avg rank), but semantic-similarity based metrics still perform second best (S-BertScore and S-BLEURT). The ranking of best metrics changes when the underlying test data with human ratings originates from human-written/gold reference outputs or system-generated outputs. This is also discussed by Scialom et al. (2021a) for simplicity transfer. This is likely due to the nature of the system-generated output used, where verbatim repetitions happen to correlate with human judgments of content preservation. Our test set (Table 4): Several metrics show low or no correlation, or even significant negative correlation for some metrics on the entire test set (BLEU) or on parts (Meteor, BertScore, Cosine).
Results on our test set show that similarity-based metrics for content preservation are unsuitable: Similarity-based metrics are deemed unsuitable for content preservation in style transfer (Section 2), and we can now support this claim by assessing the correlation of these metrics to human judgments on our test set. On our test set, similarity-based metrics show low or negative correlations with human judgment. We conclude that the high correlation scores previously reported on similarity-based metrics obtained on existing dataset are misleading. We hypothesise that misleading conclusions from prior meta-evaluations stem from the nature of the source data used: system-generated outputs happen to be very similar to source sentences lexically and semantically, whereas human-written/gold reference outputs do not, or at least rarely, contain samples where content not related to the style shift is poorly preserved or even fabricated - both limiting the cases where the metrics ability are tested.
@@
Content preservation: Content preservation metrics must be style-aware. Across all tasks and data sources, overall, the highest correlations between metric and human judgement are achieved by using LLM-as-Autorater using 70B LLaMA3-Instruct. If compute is a constraint, we recommend the LogProb method with a smaller model. Both model sizes of 3B and 1B outperform 3B and 1B LLM-as-Autorater  [^5] The LLM-as-Autorater 1b result is not reported as for over 50% of the cases, the output was not compliant with the expected output format (Figure 2). For a backbone model size of 8b, LLM-as-Autorater is recommended.
@@ 7 Conclusion
We conduct a large-scale meta-evaluation of metrics for style transfer, focusing on content preservation. We construct a new challenging test set for assessing metrics for content preservation with human judgement. We show that the meta-evaluation of metrics for content preservation is not straightforward. Previous meta-evaluation studies on existing datasets consisting of either automatically generated data from specific systems or human-generated data alike leads to misleading conclusions about best metrics, due to skewed test cases. On existing system-generated data, human judgment on content preservation happens to correlate with verbatim repetitions, and on gold references/human-written data, the limits of the metrics are not tested, as cases of drastically low preserved or fabricated content are not or rarely present in the data. With our new test set, we demonstrate empirically that similarity-based metrics are not suitable for content preservation for style transfer. Instead, metrics for content preservation should be style-aware. We propose a new efficient style-aware metric for content preservation and one for style strength, reusing the same computational effort (LogProb). Overall, we find that a large LLM-as-Autorater (70B) achieves the highest correlation between content preservation scores and human judgments. However, if considering computational efficiency or feasibility in the form of a smaller model, then our method LogProb outperforms a LLM-as-Autorater baseline using a 3B parameter LLM.
@@ Acknowledgements
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned $\end{array}$ This work was supported by the Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429) and VILLUM FONDEN (40516). It was further supported by the European Union (ERC, ExplainYourself, 101077481), and by the Pioneer Centre for AI, DNRF grant number P1.
@@ Appendix A Data sets
Appendix B Synthetic Generated Test Part

### v4

@@ Abstract
Large language models (LLMs) make it easy to rewrite a text in any style – e.g. to make it more polite, persuasive, or more positive – but evaluation thereof is not straightforward. A challenge lies in measuring content preservation: that content not attributable to style change is retained. This paper presents a large meta-evaluation of metrics for evaluating style and attribute transfer, focusing on content preservation. We find that meta-evaluation studies on existing datasets lead to misleading conclusions about the suitability of metrics for content preservation. Widely used metrics show a high correlation with human judgments despite being deemed unsuitable for the task – because they do not abstract from style changes when evaluating content preservation. We show that the overly high correlations with human judgment stem from the nature of the test data. To address this issue, we introduce a new, challenging test set specifically designed for evaluating content preservation metrics for style transfer. We construct the data by creating high variation in the content preservation. Using this dataset, we demonstrate that suitable metrics for content preservation for style transfer indeed are style-aware. To support efficient evaluation, we propose a new style-aware method that utilises small language models, obtaining a higher alignment with human judgements than prompting a model of a similar size as an autorater.
@@ 1 Introduction
Figure 1: Sample from our new test set with human annotations. S: style strength, C: content preservation. Metrics for content preservation: Bold indicates the highest rated output, and the checkmark indicates successful transfer.
@@
Our aggregated results (Figure 2) demonstrate that similarity-based metrics for content preservation show low to negative correlations with human judgment on our test set (last column), but exhibit overly high correlations to human judgment on existing data sets (first two columns). Style-aware approaches show positive correlations to human judgments on all data types (all columns). Note, our dataset is designed to stress-test the metrics, and we therefore recommend assessing all data types for evaluating the metrics.
Figure 2: Average Spearman correlation between metrics and human judgments for content preservation, grouped by whether test data originates from system-generated output (total of 5 sources), human-written (total of 3 sources) or our constructed test set. Which metric is deemed best changes depending on the test data, and only our test set correctly finds similarity-based metrics to be unsuitable.
@@ 2 Background and Related Work
Evaluation. Traditionally, style transfer systems have been evaluated along three dimensions: 1) style strength – did the style successfully shift?; 2) content preservation – is the content otherwise the same?; 3) fluency – is the rewrite fluent and grammatically correct? Jin et al. (2022). [^2] Note that fluency is not covered in this work Some work on style transfer systems includes small-scale human evaluations of the systems, but papers heavily rely on automatic metrics for evaluation Ostheimer et al. (2023):
Style. The de facto evaluation approaches, also in recent work, use either a classifier, e.g. RoBERTa, trained on a specific style Lai et al. (2024); Hallinan et al. (2023); Han et al. (2024); Mukherjee et al. (2024); Liu et al. (2024); Luo et al. (2023); Zeng et al. (2024), or a regressor Briakou et al. (2021a); Lai et al. (2022); Pauli et al. (2025). However, both require training data for the specific style, limiting their scope and scale. Further, Briakou et al. (2021a) criticises the use of binary classifiers, as this approach does not capture that a style can shift to differing degrees. In this paper, we focus on evaluating arbitrary styles without trained classifiers or regressors, and we focus on approaches that output scores to assess different degrees of style changes.
@@
A proposed solution to this criticism is to make metrics for content preservation style-aware. Mir et al. (2019) proposes to use a lexicon of style words to mask the sentence in a sentiment task. In a follow-up study, Yu et al. (2021) extract a style lexicon automatically. More recent work prompts an LLM for evaluation (LLM as an autorater, Zeng et al. (2024); Lai et al. (2023); Ostheimer et al. (2024)). For other natural language generation tasks, LLM-as-a-judge approaches are deemed better than similarity-based metrics when tasks are semantically nuanced Li et al. (2024).
Still, despite the established criticism of the use of semantic similarity metrics for content preservation in style transfer, these metrics often show reasonable correlation with human judgment on content preservation in style transfer. Prior meta-evaluation efforts (Lai et al., 2023; Ostheimer et al., 2024) show that using LLMs for evaluating content preservation is only on par rather than superior to using semantic similarity-based metrics.
@@ 3 Methodology: Test Set
Human annotators rate style shift on a 1–5 scale (5 = very good). Both outputs are seen as successfully shifting to a more polite tone, but Output 2 scores poorly on content preservation. Yet, similarity-based metrics wrongly rate Output 2 as better at preserving content than Output 1, while style-aware metrics correctly assign it a lower score. The employed metrics are formally introduced in Section 5.
@@
We aim to construct data where all output sentences succeed on style transfer, but for each source sentence, only one output preserves the content well, creating variation in content preservation quality.
Other examples from our test set include the task of rewriting the following sentence to be detoxified; yeah but she is making me feel like shit; with the two rewritings:
  • Output 1: Yes I know, but she is really making me feel terrible
  • Output 2: yeah but she is making me feel
Another example is a neutral headline from WikiNews to be framed more positively; President of China lunches with Brazilian President; with the rewritings;
  • Output 1: The Great Presidents of China and Brazilian strengthen important ties over lunchs
  • Output 2: The President of China enjoys lunches with the Brazilian first lady
In total, our constructed test data covers six style/attribute tasks and consists of 500 samples, 100 of which are manually created and the remaining generated using an LLM for scale.
Manually created. We construct 100 transfer samples from a source sentence by manually creating two output sentences as described. We create: Task 1) on sentiment (positive/negative), using headlines from Wiki News [^3] Wikinews.org as source sentences with minor modifications; Task 2) on detoxifying, using toxic sentences from Logacheva et al. (2022b). We supply each sample with a gold-truth reference sentence to enable a small evaluation of reference-based methods.
LLM generated. We use GPT-4o-mini444openai.com to create style transfer pairs in a multi-step process:
@@
We construct the LLM-generated part of the test set with four different style transfer tasks: i) a headline to be more catchy; ii) an impolite sentence from an email to be polite; iii) a persuasive request to be more persuasive; and iv) a sentence with informal language, with grammatical mistakes and internet slang to be formal. Prompt details are included in App. B, where we also include an automatic assessment of the ‘linguistic acceptability’ of the samples.
@@ 3.1 Human Annotations
We multi-annotate our samples and obtain a good level of agreement: Three workers annotate each sample in batches, grouped by style task, with a total of five different workers. Annotators rate, on a 5-point Likert scale, how well the style or attribute change is achieved, and how well the meaning/content unrelated to the style/attribute change is preserved, following previous work Mir et al. (2019); Ziegenbein et al. (2024). We use the scale
@@
We achieve a high level of human agreement on content preservation with a Krippendorff’s Alpha Krippendorff (2011) of $0.768$, and a good level of success in style transfer, with an average score on the 5-point Likert scale of $4.27$. Detailed results per task in Table 1, where we report the percentage of samples that at least obtain an average rating of 3 (fair). Most sentences successfully transfer style.
The workers are recruited via a crowdsourcing platform (prolific.com). The five different workers who participate are from the UK, hold a BA in Arts and are experienced on the platform. Workers are paid a fixed amount per batch at a rate, which the platform considers a ’great’ hourly pay. Details on annotation guideline, setup and payment are provided in App. C.
@@
catchy LLM 100 0.806 90
polite LLM 100 0.645 100
persuasive LLM 100 0.80 88
formal LLM 100 0.817 99
Table 1: Stats on our test set. ma: manually created, LLM: LLM-generated. C: content preservation, S: style strength. All samples are manually annotated by three raters, and IAA shows the inter-annotator-agreement using Krippendorff’s alpha.
@@ 4 Methodology: LogProb
We propose a novel approach for evaluating style transfer with the following properties: 1) no training data required on style; 2) no need for gold truth / reference sentences; 3) improved feasibility in the form of model size (e.g. 1B,3B) with improved performance compared to our baseline of prompting an LLM-as-Autorater. For content preservation, we additionally aim for: 4) accounting for changes in target style – ensuring theoretical suitability as discussed in Section 2. For style strength, beyond properties 1,2 and 3, we aim for: 5) scoring on a range rather than a binary score – motivated by the fact that styles can shift to various degrees as discussed in Section 2 and recommended by Briakou et al. (2021a).
@@
and $t^{s}$$t^{pa}$$t^{r}$ being instructions on rewrite with respect to style, paraphrase and repeat.
@@
$S=\frac{1}{n}\sum_{i=1}^{n}\left(p_{i}^{t^{s}}-\max\left(p_{i}^{t^{pa}},p_{i}^{t^{r}}\right)\right)$ (2)
We deploy three backbone model sizes—1B,3B and 8B parameters—using Llama 3 Instruct models Dubey et al. (2024).555 meta-llama/Llama-3.2-3B-Instruct,meta-llama/Llama-3.2-1B-Instruct and meta-llama/Llama-3.1-8B-Instruct. We deploy different instructions as part of the context as per Table 2. Further details in App. D.
@@ 5 Benchmarking
    Semantic similarity: token-based BERTScore Zhang et al. (2019); Cosine similarity between sentence embeddings Feng et al. (2022); learned BERT-based BLEURT Sellam et al. (2020). COMET Li et al. (2024), which uses both source and references to evaluate; however, COMET also have a reference-free version.
@@ 6 Results
Meta-evaluation is highly impacted by the type of test samples. Overall metric performance (correlation with human judgements) varies greatly depending on whether the human-rated outputs are system-generated, human-written, or from our newly constructed test set (Fig. 2). On system-generated data (Table 3): Many metrics show relatively high correlations across tasks. For all similarity-based metrics except COMET, the versions using source sentences [S-] result in higher correlations to human ratings than the versions using references [R-]. The best metrics overall are S-BertScore and S-BLEURT (avg+avg rank). On gold reference data (Table 3): Here, LLM-as-Autorater 70B performs best (avg+avg rank), but semantic-similarity based metrics still perform second best (S-BertScore and S-BLEURT). The ranking of best metrics changes when the underlying test data with human ratings originates from human-written/gold reference outputs or system-generated outputs. This is also discussed by Scialom et al. (2021a) for simplicity transfer. This is likely due to the nature of the system-generated output used, where verbatim repetitions happen to correlate with human judgments of content preservation. Our test set (Table 4): Several metrics show low or no correlation, or even significant negative correlation for some metrics on the entire test set (BLEU) or on parts (Meteor, BertScore, Cosine).
Results on our test set show that similarity-based metrics for content preservation are unsuitable: Similarity-based metrics are deemed unsuitable for content preservation in style transfer by prior work (Section 2), and we can now support this claim by assessing the correlation of these metrics to human judgments on our test set. On our test set, similarity-based metrics show low or negative correlations with human judgment. We conclude that the high correlation scores previously reported on similarity-based metrics obtained on existing dataset are misleading. We hypothesise that misleading conclusions from prior meta-evaluations stem from the nature of the source data used: system-generated outputs happen to be very similar to source sentences lexically and semantically, whereas human-written/gold reference outputs do not, or at least rarely, contain samples where content not related to the style shift is poorly preserved or even fabricated – both limiting the cases where the metrics ability are tested.
@@
Content preservation: Content preservation metrics must be style-aware. Across all tasks and data sources, overall, the highest correlations between metric and human judgement are achieved by using LLM-as-Autorater using 70B Llama 3 Instruct. If compute is a constraint, we recommend the LogProb method with a smaller model. Both model sizes of 3B and 1B outperform 3B and 1B LLM-as-Autorater  [^6] The LLM-as-Autorater 1b result is not reported as for over 50% of the cases, the output was not compliant with the expected output format (Figure 2). For a backbone model size of 8b, LLM-as-Autorater is recommended.
@@ 7 Conclusion
We conduct a large-scale meta-evaluation of metrics for style transfer, focusing on content preservation. We construct a new challenging test set for assessing metrics for content preservation with human judgment. We show that the meta-evaluation of metrics for content preservation is not straightforward. Previous meta-evaluation studies on existing datasets consisting of either automatically generated data from specific systems or human-generated data alike leads to misleading conclusions about best metrics, due to skewed test cases. We hypothesise that on existing system-generated data, human judgment on content preservation happens to correlate with verbatim repetitions, and on gold references/human-written data, the limits of the metrics are not tested, as cases of drastically low preserved or fabricated content are not or rarely present in the data. With our new test set, deliberately designed with variation in content preservation, we demonstrate empirically that similarity-based metrics are not suitable for content preservation for style transfer. Instead, metrics for content preservation should be style-aware. We recommend for meta-evaluation that a diverse set of data sources be used, including both system-generated test cases with human ratings, as well as our newly constructed test set, which tests the nature of content preservation metrics.
We propose a new efficient style-aware metric for content preservation and one for style strength, reusing the same computational effort (LogProb).
Overall, we find that a large LLM-as-Autorater (70B) achieves the highest correlation between content preservation scores and human judgments. However, if considering computational efficiency or feasibility in the form of a smaller model, then our method LogProb outperforms a LLM-as-Autorater baseline using a 3B parameter LLM.
@@ Acknowledgements
[…]
@@ Appendix A Data sets
Appendix B LLM-generated Test Part
@@
Automatic assessment of ‘linguistic acceptability’
To assess the quality of the LLM-generated test samples compared to the manually created ones, we assess the ‘linguistic acceptability’ of the samples. We use a pretrained binary classifier from HuggingFace.com [textattack/roberta-base-CoLA], trained on the CoLA dataset on grammatical acceptability Warstadt et al. (2019). We report the percentage of the samples predicted as ’linguistically acceptable’ grouped by LLM-generated and manually created sampels. We observe that most samples are predicted as grammatically acceptable, Table 7.
Manually created LLM-generated
source 94 % 93.5 %
rewrites 80 % 98.75 %
Table 7: Percentage of samples predicted as lingusitic acceptable